OpenStudy (anonymous):

if you a function such as f(x)= (x^2+4x-12)/(x^2-2x) if you wanna find lim x->2, you do f(2) is that right

OpenStudy (anonymous):

is this because limx->a(fx)=f(a) ??

OpenStudy (calculusfunctions):

NO!!!!!!!!!!!!!!

OpenStudy (calculusfunctions):

TO FIND A LIMIT OF A FUNCTION IS NOT THE SAME AS FINDING WHERE THE FUNCTION IS DISCONTINUOUS!!!!!

OpenStudy (anonymous):

im not finding discontinuous?? i wanna find the lim x->2

OpenStudy (calculusfunctions):

Yes I know, that is exactly what I have been trying to tell you since you've been messaging me! In the messages I already told you that since it is an indeterminate form of a rational function, you first simplify the rational function and then evaluate the limit.

OpenStudy (anonymous):

so if u get a indeterminate form you simplify and do f(a)

OpenStudy (anonymous):

@calculusfunctions WILL YOU HAVE TIME TO HELP ME WITH 2 QUESTIONS when ur done here

OpenStudy (calculusfunctions):

Example:\[\lim_{x \rightarrow -2}\frac{ 4-x ^{2} }{ 2+x }\]If you substitute x = -2 into this rational function, you'll find that f(-2) = 0/0. Thus this limit is an indeterminate form. Hence we will first simplify the rational function by factoring both the numerator and the denominator as much and if possible, then divide any like factors. Thus\[\lim_{x \rightarrow -2}\frac{ 4-x ^{2} }{ 2+x }\] \[=\lim_{x \rightarrow -2}\frac{ (2-x)(2+x) }{ 2+x }\] \[=\lim_{x \rightarrow -2}(2-x)\] \[=2-(-2)\] \[=4\]Therefore\[\lim_{x \rightarrow -2}\frac{ 4-x ^{2} }{ 2+x }=4\]

OpenStudy (anonymous):

fanks teacher. line after limx->-2(2-x) you subbed x=-2 right. why is that a valid step? is that because limx->af(x)=f(a)?